OpenStudy (vuriffy):

I really need help on a question for my test corrections, I don't have a clue. Any help is appreciated! 4cos(x) ------ + 15 tan(x) = 4sin^2(x) - 15sin(x) - 4

OpenStudy (jdoe0001):

so.. .what are you supposed to do there anyway?

OpenStudy (vuriffy):

\[4\cos(x) / \tan(x) + 15 = 4\sin^2(x) - 15\sin(x) - 4\]

OpenStudy (vuriffy):

Show that the equation (first one) can be expressed as the (second one).

OpenStudy (vuriffy):

They are both equal to zero, which is why they are equal to each other.

OpenStudy (jdoe0001):

hmmm have you covered trigonometric identities yet?

OpenStudy (vuriffy):

Yes, I can get my notebook, if you want me to. It's only for my test corrections, I am just beyond confused. Hold on a second, please.

OpenStudy (jdoe0001):

hmmm lemme check about

OpenStudy (vuriffy):

Like, I know how to do the second part of the question, where you have to find the numbers between 0 < x < 360. I just have trouble with converting them to equal each other. I have my notebook now.

OpenStudy (vuriffy):

I have the information to do it, I just don't know what to do in the case of this problem.

OpenStudy (jdoe0001):

tis \(\bf \cfrac{4cos(x)}{tan(x)}+15=4sin^2(x)-15sin(x)-4\ right?\) is there a -4 on the RHS?

OpenStudy (vuriffy):

Yes.

OpenStudy (danjs):

i put a value for x angle, evaluating it, the equality is not the same

OpenStudy (jdoe0001):

hmm lemme gimme a few seconds lemme post what I have

OpenStudy (vuriffy):

In this case, we are not finding the value of x. We are converting the one equation into the other.

OpenStudy (vuriffy):

With identities. I just do not know how to convert it correctly to equal the other one.

OpenStudy (danjs):

i picked a random test angle, the left and right sides of the equality give different values

OpenStudy (vuriffy):

Danjs, the second part of the problem has two different values, yes. Can you tell me what you got, because I have the correct values for that part.

OpenStudy (jdoe0001):

\(\bf \cfrac{4cos(x)}{tan(x)}+15=4sin^2(x)-15sin(x)-4 \\ \quad \\ \quad \\ \cfrac{4cos(x)}{tan(x)}+15\implies \cfrac{4cos(x)+15tan(x)}{tan(x)} \\ \quad \\ \cfrac{4cos(x)+15{\color{brown}{ \frac{sin(x)}{cos(x)}}}}{tan(x)}\implies \cfrac{4cos(x)+\frac{15sin(x)}{cos(x)}}{tan(x)}\implies \cfrac{\frac{4cos^2(x)+15sin(x)}{cos(x)}}{\frac{sin(x)}{cos(x)}} \\ \quad \\ \cfrac{4cos^2(x)+15sin(x)}{\cancel{cos(x)}}\cdot \cfrac{\cancel{cos(x)}}{sin(x)}\implies \cfrac{4cos^2(x)+15sin(x)}{sin(x)} \\ \quad \\ \cfrac{4[{\color{brown}{ 1-sin^2(x)}}]+15sin(x)}{sin(x)}\implies \cfrac{4-4sin^2(x)+15sin(x)}{sin(x)} \) which doesn't quite equate the RHS btw

OpenStudy (jdoe0001):

the numerator comes close, but it has the wrong sign but doesn't quite equate the RHS though

OpenStudy (vuriffy):

The issue is the sin(x) on the bottom?

OpenStudy (vuriffy):

The numerator rearranged equals the other equation, yes?

OpenStudy (jdoe0001):

anyhow.. need to dash but they're not equal

OpenStudy (vuriffy):

Mhm, I see that. Logically for the question, they are supposed to equal each other in some way, or form. I believe you have given me enough information to help me finish it myself. I am greatly pleased with your help, as an old member of Open Study, I am grateful for the help. You have an wonderful day, or night. You aswell, DanJS.